3 fl. oz of a 45% saline solution was mixed wit 7 fl. oz. of a 35% saline solution. What is the concentration of the mixture

Answer by checkley79(3341)

.45*3+.35*7=10X
1.35+2.45=10X
3.8=10X
X=3.8/10
X=.38 OR 38% IS THE MIXTURE.